# Matching Markets Meet LLMs: Algorithmic Reasoning with Ranked Preferences

### v1

@@ I.8 Example Prompt for Repeated Queries Due to Incomplete Matching
    Alright, I have been asked to determine whether agent W3 would prefer M4 and M3 over M1. For this, let me look at W3’s preference list in order: position 1: M4 position 2: M5 position 3: M1 position 4: M3 position 5: M2 So, M4 is W3’s 1-most preferred partner, M3 is W3’s 4-most preferred partner, and M1 is W3’s 3-most preferred partner. Hence, agent W3 prefers M4 over M1, and agent W3 does not prefer M3 over M1. This means that the order of preference is M4 $>$ M1 $>$ M3. Therefore, since the question was whether agent W3 would prefer BOTH M4 AND M3, given that it DOES NOT prefer M3 over M1, the answer is ’NO’. $<$/think$>$
@@
    Alright, I have been asked to determine whether agent W5 would accept a proposal from M1 or M4, if she is currently matched with M3. For this, let me look at W5’s preference list in order: position 1: M5 position 2: M4 position 3: M2 position 4: M3 position 5: M1 So, M1 is W5’s 5-most preferred partner, M4 is W5’s 2-most preferred partner, and M3 is W5’s 4-most preferred partner. Hence, agent W5 does not prefer M1 over their current partner M3, and agent W5 prefers M4 over their current partner M3. This means that the order of preference is M4 $>$ M3 $>$ M1. Therefore, since the question was whether agent W5 would prefer EITHER M1 OR M4, given that it prefers M4 over M3, the answer is ’YES’. $<$/think$>$
J.6 Answer ($\boldsymbol{a^{(i)}}$):

### v2

@@ Appendix I Inference Details
    Alright, I have been asked to determine whether agent W3 would prefer M4 and M3 over M1. For this, let me look at W3’s preference list in order: position 1: M4 position 2: M5 position 3: M1 position 4: M3 position 5: M2 So, M4 is W3’s 1-most preferred partner, M3 is W3’s 4-most preferred partner, and M1 is W3’s 3-most preferred partner. Hence, agent W3 prefers M4 over M1, and agent W3 does not prefer M3 over M1. This means that the order of preference is M4 >> M1 >> M3. Therefore, since the question was whether agent W3 would prefer BOTH M4 AND M3, given that it DOES NOT prefer M3 over M1, the answer is ’NO’. <</think>>
@@
    Alright, I have been asked to determine whether agent W5 would accept a proposal from M1 or M4, if she is currently matched with M3. For this, let me look at W5’s preference list in order: position 1: M5 position 2: M4 position 3: M2 position 4: M3 position 5: M1 So, M1 is W5’s 5-most preferred partner, M4 is W5’s 2-most preferred partner, and M3 is W5’s 4-most preferred partner. Hence, agent W5 does not prefer M1 over their current partner M3, and agent W5 prefers M4 over their current partner M3. This means that the order of preference is M4 >> M3 >> M1. Therefore, since the question was whether agent W5 would prefer EITHER M1 OR M4, given that it prefers M4 over M3, the answer is ’YES’. <</think>>
K.6 Answer ($\boldsymbol{a^{(i)}}$):